Question 4 (of 5) 10.00 points Find the sample proportions and test statistic for equal proportions (a-1) Dissatisfied workers in two companies 저 Ⅱ 38, n1 100, x2 :28, n2 : 100, a = 05, two-tailed test Round your answers to 4 decimal places. Use Excel to calculate the p-value.) pt P2 p value 702 (a-2) Choose the appropriate hypotheses a 3) Based on the dala soject Ho True False 0-14-4xs AStat 1 Exam 1 Re.docx 4 Stat il Exam 1 Re..docx Type here to search 46 3 4 5 6 7 8 9
(b-1) Rooms rented at least a week in advance at two hotels x1 = 22, n17 200, x2 = 10, n2 = 50, a-01, ett-tailed test. (Round your answers to 4 decimal places. Negative values should be indicated by a minus sign. Use Excel to calculate the p-value.) P1 P2 Zcale р.value za (b-21 Choose the appropriate hypotheses b-2) Based on the data reject H True False (c-1) Horne eq ity loan default rates in two banks r,-34, n: 480. x2 : 24. n2: 520 о» 05, righeta led test (Round your answers to 4 decimal places. Use Excel to calculate the p-value.) Pr Stat Il Exam 1 Re...doc Stat II Exam 1 Re...docx a Stat Exam 1 Re.. docx lype here to search 4 5
Answer #1

Part a)

Given,

x1 = 38

n1 = 100

x2 = 28

n2 = 100

alpha = 0.05

a1)

p1 = 38/100 = 0.38

p2 = 28/100 = 0.28

pc = (n1p1 + n2p2)/(n1+n2) = 0.33

ZCalc = (p1-p2)/sqrt(pc * (1-pc)/n1 + pc * (1-pc)/n2) = 1.5038

p-value = .132633

Zalpha/2 = +/-1.96

a2)

Option a

a3)

Option b ie False

Part b)

Given,

x1 = 22

n1 = 200

x2 = 10

n2 = 50

alpha = 0.01

b1)

p1 = 22/200 = 0.11

p2 = 10/50 = 0.2

pc = (n1p1 + n2p2)/(n1+n2) = 0.128

ZCalc = (p1-p2)/sqrt(pc * (1-pc)/n1 + pc * (1-pc)/n2) = -1.7038

p-value = .044284

Zalpha = -2.326

b2)

Option b

b3)

Option b ie False

Part c)

Given,

x1 = 34

n1 = 480

x2 = 24

n2 = 520

alpha = 0.05

c1)

p1 = 34/480 = 0.0708

p2 = 24/520 = 0.0462

pc = (n1p1 + n2p2)/(n1+n2) = 0.058

ZCalc = (p1-p2)/sqrt(pc * (1-pc)/n1 + pc * (1-pc)/n2) = 1.6681

p-value = .047648

Zalpha = 1.96